# Optimality of quasi-Monte Carlo methods and suboptimality of the sparse-grid Gauss–Hermite rule

### v1

@@ 4.2 QMC with Möbius transformation
$=\sum_{j=0}^{r-\eta+1}\binom{r}{j}\Phi^{(j+1)}(t)\;B_{r-j,\eta-1}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-j-\eta+2)}(t)\bigr).$
They also satisfy another relation
$\partial_{j}B_{r,\eta}(x_{1},\ldots,x_{r-\eta+1})$ $=\binom{r}{j}B_{r-j,\eta-1}(x_{1},\ldots,x_{r-j-(\eta-1)+1}),\quad j=1,\dots,r-\eta+1,$
and thus differentiating $C_{r,\eta}$ gives
@@
Hence, subtracting $C_{r+1,\eta}(t)$ from $C^{\prime}_{r,\eta}(t)$ completes the proof. ∎
@@
$C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t}):=\prod_{j=1}^{d}C_{j;r_{j},\eta_{j}}(t_{j}).$
@@
$D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ (18)
@@
We now re-index the first sum so that the summation over $\boldsymbol{\eta}$ on the right-hand side runs up to $\mathbf{r}+\mathbf{e}_{j}$. Since only the $j$-th component of $\boldsymbol{\Phi}$ depends on $t_{j}$, we have $\partial_{{}_{j}}\bigl(v_{\boldsymbol{\eta}}\circ\boldsymbol{\Phi}\bigr)(\boldsymbol{t})=v_{\boldsymbol{\eta}+\mathbf{e}_{j}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\Phi_{j}^{\prime}(t_{j})$. Moreover, $\partial_{j}C_{\mathbf{r},\boldsymbol{\eta}}=\bigl(\partial_{j}C_{j;r_{j},\eta_{j}}\bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}=0$ if $\eta_{j}=0$. Combining these, we obtain
@@
$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})$ $=\Phi_{j}^{\prime}(t_{j})C_{j;r_{j},r_{j}}(t_{j})\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})$
$=C_{j;r_{j}+1,r_{j}+1}(t_{j})\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t})$
@@
$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ $=\Bigl(\Phi_{j}^{\prime}(t_{j})C_{j;r_{j},\eta_{j}-1}(t_{j})+\partial_{j}C_{j;r_{j},\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})$
$=\Bigl(C_{j;r_{j}+1,\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{k;r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t}).$
@@
which is (18) for $\mathbf{r}+\mathbf{e}_{j}$. This completes the proof. ∎
@@
$D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ (19)
@@
Local integrability follows from the boundedness of $(\Phi_{j}^{-1})^{\prime}$ and the derivatives $\Phi_{j}^{(\tau)}$ on compact subsets of $(0,1)^{d}$, as well as from the facts that $D^{\boldsymbol{\eta}}v$ is locally integrable on $\mathbb{R}^{d}$, and that the image of each compact subset of $(0,1)^{d}$ under $\boldsymbol{\Phi}$ is compact in $\mathbb{R}^{d}$. To show the equality above, fix $\varphi$ and take an open set $U$ such that $\mathrm{supp}(\varphi)\subset U\Subset(0,1)^{d}$. Noting that $(\Phi_{j}^{-1})^{\prime}$ and $\Phi_{j}^{(\tau)}$ are bounded on $U$, we invoke Lemma 4.5 together with a standard argument for change-of-variables for Sobolev functions to conclude that the equality holds on $U$, and thus equivalently on $(0,1)^{d}$. ∎
Note that, unlike in the common setting for Sobolev spaces, we do not assume boundedness of the derivatives of $\boldsymbol{\Phi}$, since the derivatives of the change of variables $\boldsymbol{\Phi}_{\mathrm{cot}}$ we consider are unbounded. As a result, and in contrast to standard results, the integrability of $D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})$ does not follow immediately, and the membership of $(f\cdot\rho)\circ\boldsymbol{\Phi}_{\mathrm{cot}}$ in the Sobolev space is handled separately in the next proposition.
@@
$\|D^{\mathbf{r}}g\|_{L^{2}([0,1]^{d})}\leq C\|f\|_{H_{\rho}^{\alpha}},$ (20)
where $C>0$ is a constant independent of $\mathbf{r}$ and $f$; for every $\ell=1,\dots,d$, every $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$ such that $r_{\ell}\leq\alpha-1$, and almost every $(t_{1},\dots,t_{\ell-1},t_{\ell+1},\dots,t_{d})\in(0,1)^{d-1}$,
$\lim_{t_{\ell}\to 0+}D^{\mathbf{r}}g(\boldsymbol{t})=\lim_{t_{\ell}\to 1-}D^{\mathbf{r}}g(\boldsymbol{t})=0,$ (21)
@@
$\leq\bigg|\Big(([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big((\rho^{1/2+\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\bigg|$ (22)
@@
so $\bigl\|D^{\mathbf{k}}\bigl(\rho^{1/2+\varepsilon}D^{\boldsymbol{\lambda}_{1}}f\bigr)\bigr\|_{L^{2}(\mathbb{R}^{d})}\leq\sum_{\mathbf{j\leq k}}\binom{\mathbf{k}}{\mathbf{j}}c_{\mathbf{j},\varepsilon}\|\rho^{1/2}D^{\mathbf{k}-\mathbf{j}+\boldsymbol{\lambda}_{1}}f\|_{L^{2}(\mathbb{R}^{d})}<\infty$. Showing $v\in W_{\mathrm{mix}}^{1,2}(\mathbb{R})$ when we only have $r_{\ell}\leq\alpha-1$ is analogous. As a result, noting that $(\Psi_{\mathrm{cot}}^{-1},\dots,\Psi_{\mathrm{cot}}^{-1})\colon\mathbb{R}^{d-1}\to(0,1)^{d-1}$ maps null sets to null sets, we see that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})$ admits a continuous representative $(0,1)^{d-1}$-a.e. For $|\mathbf{r}|_{\infty}\leq\alpha-1$, the function $\rho^{1/2+\varepsilon}\cdot D^{\boldsymbol{\lambda}_{1}}f$ is in $W_{\text{mix}}^{1,2}(\mathbb{R}^{d})$ and thus the analogous claim holds everywhere in $(0,1)^{d-1}$.
@@
where in the last line, we again used the fact that $([D^{\tau}\phi]\circ\phi^{-1}(x_{k})\big)^{\ell_{k,\tau}}$ is a polynomial of degree $(\tau+1)\ell_{k,\tau}$; see Lemma 4.3. Thus we have proved (21). ∎
@@
For the Korobov space, the rank-$1$ lattice rule constructed by [12, Algorithm 3.14] achieves the convergence rate $N^{-\alpha}(\ln N)^{d\alpha}$. For the unanchored Sobolev space, higher-order digital nets achieve the optimal rate $N^{-\alpha}(\ln N)^{(d-1)/2}$ [29]. This rate is optimal, including the logarithmic factor, among all linear and nonlinear quadrature classes that rely solely on function evaluations. QMC rules for $\mathbb{R}^{d}$ with Möbius transformation inherit these results: those based on rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas those based on higher-order digital nets inherit full optimality, including the logarithmic factor, as we show below.
Remark 1 (Evaluation at the boundary).
@@
Let $Q_{N}$ be the rank-$1$ lattice rule with cotangent transform as in (17). Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [12, Algorithm 3.14], $Q_{N}$ satisfies
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{d\alpha}.$ (23)
Proof.
For $f\in H_{\rho}^{\alpha}$, set $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$ for $\boldsymbol{t}\in(0,1)^{d}$. Fix $\ell\in\{1,\dots,d\}$ and choose $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$, $r_{\ell}\leq\alpha-1$. Noting that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}$ is absolutely continuous on the interval $(0,1)$ for a.e. $\boldsymbol{t}_{-\ell}\in(0,1)^{d}$, we use integration by parts in $(0,1)$ [44, Corollary 3.23], as well as uniformly continuity up to $[0,1]$ and the integrability of functions under consideration on $(0,1)$, to obtain
@@
Finally, since $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C\|g\|_{\mathrm{Kor},\alpha}N^{-\alpha}(\ln N)^{d\alpha}$, the statement follows. ∎
@@
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{(d-1)/2}.$ (24)
Proof.
From [29, Theorem 1], the QMC under study $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C_{d,\alpha}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}N^{-\alpha}(\ln N)^{(d-1)/2}$, where $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$, and $\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}$ is defined in (12). From Proposition 4.7, this norm can be bounded in terms of $\|f\|_{H_{\rho}^{\alpha}}$ as
$\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}^{2}\leq\sum_{\mathbf{r}\in\{0,\dots,\alpha\}^{d}}\|D^{\mathbf{r}}g\|_{L^{2}((0,1)^{d})}^{2}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}}.$
This completes the proof. ∎
We summarise the results thus far in the following theorem.
Theorem 4.10.
The quasi-Monte Carlo rule $Q_{N}$ (9) achieves the error bound
where the exponent $\gamma$ depends on the QMC construction:
[…]

### v2

@@ 4.2 QMC with Möbius transformation
$=\sum_{{q}=0}^{r-\eta+1}\binom{r}{{q}}\Phi^{({q}+1)}(t)\;B_{r-{q},\eta-1}\bigl(\Phi^{\prime}(t),\ldots,\Phi^{(r-{q}-\eta+2)}(t)\bigr).$
and thus, together with Lemma 4.7, differentiating $C_{r,\eta}$ gives
@@
Hence, subtracting $C_{r+1,\eta}(t)$ from $C^{\prime}_{r,\eta}(t)$ completes the proof.
@@
$C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t}):=\prod_{j=1}^{d}C_{r_{j},\eta_{j}}(t_{j}).$
@@
(19) $D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$
@@
We now re-index the first sum so that the summation over $\boldsymbol{\eta}$ on the right-hand side runs up to $\mathbf{r}+\mathbf{e}_{j}$. Since only the $j$-th component of $\boldsymbol{\Phi}$ depends on $t_{j}$, we have $\partial_{{}_{j}}\bigl(v_{\boldsymbol{\eta}}\circ\boldsymbol{\Phi}\bigr)(\boldsymbol{t})=v_{\boldsymbol{\eta}+\mathbf{e}_{j}}(\boldsymbol{\Phi}(\boldsymbol{t}))\,\Phi_{j}^{\prime}(t_{j})$. Moreover, $\partial_{j}C_{\mathbf{r},\boldsymbol{\eta}}=\bigl(\partial_{j}C_{r_{j},\eta_{j}}\bigr)\prod_{k\neq j}C_{r_{k},\eta_{k}}=0$ if $\eta_{j}=0$. Combining these, we obtain
@@
$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})$ $=\Phi_{j}^{\prime}(t_{j})C_{r_{j},r_{j}}(t_{j})\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})$
$=C_{r_{j}+1,r_{j}+1}(t_{j})\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t})$
@@
$\Phi_{j}^{\prime}(t_{j})C_{\mathbf{r},\boldsymbol{\eta}-\mathbf{e}_{j}}(\boldsymbol{t})+\partial_{{}_{j}}C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$ $=\Bigl(\Phi_{j}^{\prime}(t_{j})C_{r_{j},\eta_{j}-1}(t_{j})+\partial_{j}C_{r_{j},\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})$
$=\Bigl(C_{r_{j}+1,\eta_{j}}(t_{j})\Bigr)\prod_{k\neq j}C_{r_{k},\eta_{k}}(t_{k})=C_{\mathbf{r}+\mathbf{e}_{j},\boldsymbol{\eta}}(\boldsymbol{t}).$
@@
which is (19) for $\mathbf{r}+\mathbf{e}_{j}$. This completes the proof.
@@
(20) $D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})(\boldsymbol{t})=\sum_{\boldsymbol{0}\leq\boldsymbol{\eta}\leq\mathbf{r}}(D^{\boldsymbol{\eta}}v\circ\boldsymbol{\Phi})(\boldsymbol{t})\;C_{\mathbf{r},\boldsymbol{\eta}}(\boldsymbol{t})$
@@
Local integrability follows from the boundedness of $(\Phi_{j}^{-1})^{\prime}$ and the derivatives $\Phi_{j}^{(\tau)}$ on compact subsets of $(0,1)^{d}$, as well as from the facts that $D^{\boldsymbol{\eta}}v$ is locally integrable on $\mathbb{R}^{d}$, and that the image of each compact subset of $(0,1)^{d}$ under $\boldsymbol{\Phi}$ is compact in $\mathbb{R}^{d}$. To show the equality above, fix $\varphi$ and take an open set $U$ such that $\mathrm{supp}(\varphi)\subset U\Subset(0,1)^{d}$, i.e., the closure $\overline{U}$ of $U\subset\mathbb{R}^{d}$ is compact and $\overline{U}\subset(0,1)^{d}$. Noting that $(\Phi_{j}^{-1})^{\prime}$ and $\Phi_{j}^{(\tau)}$ are bounded on $U$, we invoke Lemma 4.11 together with a standard argument for change-of-variables for Sobolev functions to conclude that the equality holds on $U$, and thus equivalently on $(0,1)^{d}$.
Having established these identities for derivatives, we are now in a position to derive bounds for the transformed integrand and its derivatives. Note that, unlike in the common setting for Sobolev spaces, we do not assume boundedness of the derivatives of $\boldsymbol{\Phi}$, since the derivatives of the change of variables $\boldsymbol{\Phi}_{\mathrm{cot}}$ we consider are unbounded. As a result, and in contrast to standard results, the integrability of $D^{\mathbf{r}}(v\circ\boldsymbol{\Phi})$ does not follow immediately, and the membership of $(f\cdot\rho)\circ\boldsymbol{\Phi}_{\mathrm{cot}}$ in the Sobolev space is handled separately in the next proposition.
@@
(21) $\|D^{\mathbf{r}}g\|_{L^{2}([0,1]^{d})}\leq C{{}_{d.\alpha}}\|f\|_{H_{\rho}^{\alpha}},$
where $C{{}_{d.\alpha}}>0$ is a constant independent of $\mathbf{r}$ and $f$ but depends on $d$ and $\alpha$; for every $\ell=1,\dots,d$, every $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$ such that $r_{\ell}\leq\alpha-1$, and almost every $(t_{1},\dots,t_{\ell-1},t_{\ell+1},\dots,t_{d})\in(0,1)^{d-1}$,
(22) $\lim_{t_{\ell}\to 0+}D^{\mathbf{r}}g(\boldsymbol{t})=\lim_{t_{\ell}\to 1-}D^{\mathbf{r}}g(\boldsymbol{t})=0,$
@@
(23) $\leq\bigg|\Big(([D^{\boldsymbol{\lambda}_{1}}f]\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\Big((\rho^{1/2+\varepsilon}\circ\boldsymbol{\Psi}_{\mathrm{cot}})(\boldsymbol{t})\Big)\bigg|$
@@
so $\bigl\|D^{\mathbf{k}}\bigl(\rho^{1/2+\varepsilon}D^{\boldsymbol{\lambda}_{1}}f\bigr)\bigr\|_{L^{2}(\mathbb{R}^{d})}\leq\sum_{\mathbf{j\leq k}}\binom{\mathbf{k}}{\mathbf{j}}c_{\mathbf{j},\varepsilon}\|\rho^{1/2}D^{\mathbf{k}-\mathbf{j}+\boldsymbol{\lambda}_{1}}f\|_{L^{2}(\mathbb{R}^{d})}<\infty$. Showing $v\in W_{\mathrm{mix}}^{1,2}(\mathbb{R})$ a.e. in $\mathbb{R}^{d-1}$ when we only have $r_{\ell}\leq\alpha-1$ is analogous. As a result, noting that $(\Psi_{\mathrm{cot}}^{-1},\dots,\Psi_{\mathrm{cot}}^{-1})\colon\mathbb{R}^{d-1}\to(0,1)^{d-1}$ maps null sets to null sets, we see that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})$ admits a continuous representative $(0,1)^{d-1}$-a.e. For $|\mathbf{r}|_{\infty}\leq\alpha-1$, the function $\rho^{1/2+\varepsilon}\cdot D^{\boldsymbol{\lambda}_{1}}f$ is in $W_{\text{mix}}^{1,2}(\mathbb{R}^{d})$ and thus the analogous claim holds everywhere in $(0,1)^{d-1}$.
@@
where in the last line, we again used the fact that $([D^{\tau}\phi]\circ\phi^{-1}(x_{k})\big)^{\ell_{k,\tau}}$ is a polynomial of degree $(\tau+1)\ell_{k,\tau}$; see Lemma 4.5. Thus we have proved (22).
@@
For the Korobov space, the rank-$1$ lattice rule constructed by [undefl, Algorithm 3.14] achieves the convergence rate $N^{-\alpha}(\ln N)^{d\alpha}$. For the unanchored Sobolev space, higher-order digital nets achieve the optimal rate $N^{-\alpha}(\ln N)^{(d-1)/2}$ [undefac]. This rate is optimal, including the logarithmic factor, among all linear and nonlinear quadrature classes that rely solely on function evaluations. QMC rules for $\mathbb{R}^{d}$ with Möbius transformation inherit these results: those based on rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas those based on higher-order digital nets inherit full optimality, including the logarithmic factor, as we show below.
Remark 4.17 (Evaluation at the boundary).
@@
Let $Q_{N}$ be the rank-$1$ lattice rule with cotangent transform as in (17). Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [undefl, Algorithm 3.14], $Q_{N}$ satisfies
(24) $\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{d\alpha}.$
Proof 4.19.
For $f\in H_{\rho}^{\alpha}$, set $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$ for $\boldsymbol{t}\in(0,1)^{d}$. Fix $\ell\in\{1,\dots,d\}$ and choose $\mathbf{r}\in\{0,\ldots,\alpha\}^{d}$, $r_{\ell}\leq\alpha-1$. Noting that $t_{\ell}\mapsto D^{\mathbf{r}}g(\boldsymbol{t})\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{t}\cdot\mathbf{k}}$ is absolutely continuous on the interval $(0,1)$ for a.e. $\boldsymbol{t}_{-\ell}\in(0,1)^{d}$, we use integration by parts in $(0,1)$ [undefat, Corollary 3.23], as well as uniformly continuity up to $[0,1]$ and the integrability of functions under consideration on $(0,1)$, to obtain
@@
Finally, since $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C\|g\|_{\mathrm{Kor},\alpha}N^{-\alpha}(\ln N)^{d\alpha}$, the statement follows.
@@
(25) $\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{(d-1)/2}.$
Proof 4.21.
From [undefac, Theorem 1], the QMC under study $Q_{N}$ satisfies $|I(f)-Q_{N}(f)|\leq C_{d,\alpha}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}N^{-\alpha}(\ln N)^{(d-1)/2}$, where $g(\boldsymbol{t})=f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\right)\prod_{k=1}^{d}\phi^{\prime}(t_{k})$, and $\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}$ is defined in (12). From Proposition 4.15, this norm can be bounded in terms of $\|f\|_{H_{\rho}^{\alpha}}$ as
$\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{0},\boldsymbol{1}])}^{2}\leq\sum_{\mathbf{r}\in\{0,\dots,\alpha\}^{d}}\|D^{\mathbf{r}}g\|_{L^{2}((0,1)^{d})}^{2}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}}^{{2}}.$
This completes the proof.